Two canoeists paddle 10km upstream in 5 hours. They make the return trip in only 1 hour. Assuming that the canoeists paddle at a constant rate, what is the speed of the current?
Found 2 solutions by TimothyLamb, lwsshak3:
Answer by TimothyLamb(4379)

x = speed of canoe in still water
y = speed of current
---
s = d/t
---
upstream:
x - y = 10/5
x - y = 2
---
downstream:
x + y = 10/1
x + y = 10
---
put the system of linear equations into standard form
---
x - y = 2
x + y = 10
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = speed of canoe in still water = 6 kph
y = speed of current = 4 kph
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by lwsshak3(11628)

Two canoeists paddle 10km upstream in 5 hours. They make the return trip in only 1 hour. Assuming that the canoeists paddle at a constant rate, what is the speed of the current?
***
let x=speed of canoe in still water
let c=speed of current
(x+c) speed of canoe downstream
(x-c) speed of canoe upstream
travel time*speed=distance
..
5(x-c)=10
1(x+c)=10
..
5x-5c=10
x+c=10
..
5x-5c=10
5x+5c=50 (mult. eq. by 5)
subtract
10c=40
c=4
speed of current=4 km/hr